tfells2004@yahoo.com
/ for fractions
Solve the equation please.
8/5r - 23/6(4/3r + 29/6) = 317/20 - 5/6r
Thank you.
Add 5/6r to both sides.
gives 48/30r +25/6r=73/30r
Multiply -23/6 *4/3r +-23/6*29/6
Solve equation for r.
Thanks again.

Answer by Alan3354(69443)

8/5r - 23/6(4/3r + 29/6) = 317/20 - 5/6r
Multiply by 60
96r - 230(4r/3 + 29/6) = 951 - 50r
Add 50r
146r - 230(4r/3 + 29/6) = 951
Multiply by 6
876r - 230(8r + 29) = 5706
876r - 1840r - 6670 = 34236
-964r = 40906
r = -20453/482